OpenStudy (anonymous):

A company will need $30,000 in 5 years. to meet this goal they will need to deposit money in an account that pays 7% annual interest compound quarterly. find the amount that will be invested to total $30,000 in 5 years.

OpenStudy (anonymous):

isnt the compound interest with the formula: \[ A=P\left(1+{r\over100k}\right)^{nk} \] r = annual rate in % k = number of times compounded per year n = number of years A = amount after all the n years P = principal amount

OpenStudy (anonymous):

we are asked to find the "P"

OpenStudy (anonymous):

ok i was using the A=P(1+r)^n

OpenStudy (anonymous):

you can use that when the principal is compounded "yearly" i.e., once a year but in this problem, it is compounded quarterly.. i.,e., 4 times a year

OpenStudy (anonymous):

oh maybe thats why i cant get it lol

OpenStudy (anonymous):

so i had 30,000=P(1+0.02)^20 is that right?

OpenStudy (anonymous):

the paranthesis is incorrect

OpenStudy (anonymous):

ok im lost..what part of it is correct?

OpenStudy (anonymous):

show me your equation before simplifying.

OpenStudy (anonymous):

\[ 30000=P\left(1+{7\over 100\times4}\right)^{5\times4} \]

OpenStudy (anonymous):

sorry i was trying to work it out. so what you just did isnt like what I did. so that would be 30000=P(1+7/400)^20?

OpenStudy (anonymous):

I just plugged in the numbers in the formula I showed in the first post

OpenStudy (anonymous):

kapeesh?